# SciTopic: Enhancing Topic Discovery in Scientific Literature through Advanced LLM

### v1

@@ Abstract
Topic discovery in scientific literature provides valuable insights for researchers to identify emerging trends and explore new avenues for investigation, facilitating easier scientific information retrieval. Many machine learning methods, particularly deep embedding techniques, have been applied to discover research topics. However, most existing topic discovery methods rely on word embedding to capture the semantics and lack a comprehensive understanding of scientific publications, struggling with complex, high-dimensional text relationships. Inspired by the exceptional comprehension of textual information by large language models (LLMs), we propose an advanced topic discovery method enhanced by LLMs to improve scientific topic identification, namely SciTopic. Specifically, we first build a textual encoder to capture the content from scientific publications, including metadata, title, and abstract. Next, we construct a space optimization module that integrates entropy-based sampling and triplet tasks guided by LLMs, enhancing the focus on thematic relevance and contextual intricacies between ambiguous instances. Then, we propose to fine-tune the textual encoder based on the guidance from the LLMs by optimizing the contrastive loss of the triplets, forcing the text encoder to better discriminate instances of different topics. Finally, extensive experiments conducted on three real-world datasets of scientific publications demonstrate that SciTopic outperforms the state-of-the-art (SOTA) scientific topic discovery methods, enabling researchers to gain deeper and faster insights [^1] The source code can be accessed through the following link: https://anonymous.4open.science/r/SciTopic-8DCC.
Index Terms:
scientific topic discovery, text clustering, large language models, document embeddings
@@ I Introduction
Recent advancements in machine learning, particularly in the realm of deep learning, have led to the emergence of various techniques aimed at automating the topic discovery process. The classical topic modeling techniques, including Latent Dirichlet Allocation (LDA) [1], Non-negative Matrix Factorization (NMF) [2], and Probabilistic Latent Semantic Analysis (PLSA) [3], could be applied to discover the scientific topics directly. However, these bag-of-words methods ignore contextual word interrelations, failing to capture the intricate semantics of modern scientific texts. In addition, these techniques often necessitate dimensionality reduction processes such as Principal Component Analysis (PCA) [4] or Uniform Manifold Approximation and Projection (UMAP) [5], potentially leading to significant information loss that is vital for maintaining the thematic depth of the documents. Deep embedding methods have gained prominence for their ability to represent textual data in a high-dimensional space, capturing semantic relationships between words and phrases. Unlike traditional bag-of-words approaches that treat words in isolation, document embeddings encapsulate the overall significance of a document in a continuous vector space, aligning semantically akin words more closely [6]. Even advanced deep topic modeling like the Embedded Topic Model (ETM) [7] and Neural Variational Document Model (NVDM) [8], though incorporating word embeddings to capture semantic nuances, still results in a limited understanding of the intricate relationships. These limitations ultimately impact the quality of insights derived from topic discovery, potentially leading to incomplete or inaccurate representations of the underlying thematic structures within the literature.
@@
To overcome these limitations, we draw inspiration from the remarkable capabilities of large language models (LLMs) in comprehending textual information. LLMs, such as GPT-4 and BERT, have exhibited unparalleled proficiency in natural language comprehension by recognizing deep contextual connections between words, phrases, and entire texts. Transformer-based architectures can capture long-range dependencies and process substantial volumes of text in a context-sensitive manner [9], effectively discerning nuanced thematic structures. Trained on vast data, transformer-based models can capture complex patterns, contextual cues, and semantic relationships beyond mere word co-occurrence. These embeddings retain syntactic and semantic linkages, offering comprehensive depictions of documents’ thematic content compared to earlier sparse models. This enhancement allows for refined topic modeling and improved semantic clustering, naturally grouping documents based on their intrinsic meanings rather than mere word frequency. For instance, Sentence-BERT (SBERT) optimizes BERT embeddings for semantic similarity tasks using a siamese network structure, preserving intricate semantic details and capturing subtle thematic distinctions [10].
@@
Along this line, by leveraging the strengths of LLMs, we propose SciTopic, an effective method to enhance scientific topic identification and provide deeper insights for researchers. Firstly, we construct a text encoder that captures essential content from scientific publications, including metadata, titles, and abstracts. Through this module, we can extract meaningful textual features crucial for accurate topic identification. Then, we introduce an LLM-guided clustering technique that leverages entropy-based sampling and triplet tasks. This approach diverges from traditional unsupervised clustering methods by actively involving the LLM in the clustering process. We utilize an entropy-based sampling strategy to identify the most ambiguous or uncertain documents, where cluster membership is less defined. These high-entropy instances are used as anchors for the triplet tasks, where two candidate titles or abstracts from nearby clusters are selected. By analyzing these triplets, the LLM refines document embeddings, sharpening the distinctions between closely related clusters. This method not only improves clustering precision but also minimizes computational overhead by focusing the LLM’s attention toward the most informative cases. Ultimately, this approach leads to a more contextually accurate and thematically coherent clustering result, ensuring that even subtle topic differences are effectively captured.
Our key contributions are as follows:
@@ II-A Neural Topic Models
Neural Topic Models (NTMs) integrate deep learning with probabilistic modeling to improve topic coherence and document representation [2, 11, 7, 12, 13, 14]. LDA inspires many neural adaptations addressing scalability and coherence for large vocabularies [1]. Recent advancements include ECRTM, which prevents topic collapse through embedding clustering regularization [15], and FASTopic, utilizing dual semantic-relation reconstruction to regulate embeddings [16]. InfoCTM aligns topics cross-lingually using mutual information [17], while BERTopic leverages pre-trained transformers and clustering for coherent topics [18]. Despite these advancements, most NTMs are not tailored for scientific topic discovery tasks.
@@ II-B Scientific Topic Discovery
Scientific topic discovery automates the identification of research trends and emerging areas, essential given the rapid growth in publications [19, 20, 21]. Domain-specific models assess research impact [22, 23]. For instance,  [24] adapts hierarchical topic discovery for scientific literature, and  [25] tracks authors’ research evolution.  [26] proposes dataset recommendation at the topic level, while  [27] identifies emerging topics by analyzing rare synonymous biterms. However, most methods lack the ability to leverage LLMs for understanding complex relationships.
@@ II-C Prompt Learning
Prompt learning advances LLM applications by designing task-specific cues, enabling few-shot and zero-shot scenarios [28, 29]. Techniques like domain-controlled prompts for remote sensing [30], graph prompt learning [31], and Match-Prompt frameworks for diverse tasks [32] have demonstrated effectiveness. Methods such as region-based image recognition [33], unsupervised image enhancement [34], and continual learning [35] further showcase prompt learning’s versatility. Inspired by prompt-based learning, we use prompts to improve query triplet generation for scientific topic discovery.
@@ III-A Problem Definition
In our task, the objective is to extract meaningful research topics from large collections of conference papers. Each paper is in the form of a document containing title ($t$), abstract ($a$), and metadata ($m$) and can represented by a set of textual features $\{x_{1},x_{2},...,x_{n}\}$}. By leveraging these features, we cluster documents into distinct groups $\{C_{1},C_{2},...,C_{k}\}$}. From each cluster, we further extract a list of key terms to verbalize the represented topic, enabling efficient topic identification and trend analysis.
@@ III-B Framework Overview
As depicted in Figure 1, our method starts with encoding the title, abstract, and metadata of each document and concatenating their embeddings to a composite feature matrix. Then we group these documents into clusters and sample ambiguous instances to construct the triplet task, which prompts the LLM to evaluate and reassign documents to more coherent clusters. These refined triplets are finally used with contrastive learning [36] to fine-tune the embedding model for an enhanced clustering performance.
@@ III-C Textual Encoder
Each document is divided into three key components: the title, abstract, and metadata. The title and abstract are encoded separately using a fine-tuned model capable of handling varied text granularities. Metadata is consolidated into a single string and embedded using the same model to maintain consistency. The encoding process is expressed:
@@
III-D LLM Guided Clustering
This section outlines a method to refine text clustering using a triplet task guided by large language models (LLMs), designed to reflect user-specific perspectives. Each triplet comprises an anchor and two candidate elements $(a,c_{1},c_{2})$, aiming to identify which candidate is more thematically similar to the anchor.
@@ III-E Fine-tuning
The fine-tuning process employs a cross-entropy loss with in-batch negative sampling [37]:
@@
where $f(x)$ is the embedding function, and $\gamma$ is a margin parameter that helps avoid trivial solutions by enforcing a minimal semantic distinction between positive and negative pairs.
@@ III-F Topic Verbalization
As depicted in Figure 2, we use class-based term frequency-inverse document frequency (c-TF-IDF) [18], which calculates scores for each cluster by treating it as a single document, emphasizing key terms that define each topic. The weight of a term $t$ in cluster $c$ is calculated as:
@@ IV-A Datasets
This study identifies prevalent topics in scholarly articles using the DBLP database and two additional Kaggle datasets. The AI-DM Research Literature Dataset, based on the DBLP database, includes metadata for 57,320 AI papers and 20,700 DM papers from major conferences such as AAAI, ACL, CVPR, IJCAI, NeurIPS, and SIGKDD, covering titles, authors, publication years, and venues. Additionally, the DBLP V10222 https://www.kaggle.com/datasets/nechbamohammed/research-papers-dataset Dataset, derived from DBLP (Version 10), spans fields like computer science, mathematics, and physics, offering metadata for 999,064 papers, including 827,533 with both titles and abstracts, from which 100,000 papers were randomly sampled. The NeurIPS333 https://www.kaggle.com/datasets/benhamner/nips-papers Dataset provides detailed information on 7,241 NeurIPS conference papers from 1987 to 2016, including titles, authors, abstracts, and full texts, offering comprehensive insights into machine learning advancements. This combination of datasets ensures robust coverage of both general and domain-specific scholarly topics for analysis. Furthermore, this study incorporates additional datasets for broader coverage, including 20,000 randomly sampled papers from the arXiv444 https://www.kaggle.com/datasets/Cornell-University/arxiv, and the PubMed Dataset555 https://www.kaggle.com/datasets/nabarupghosh/pubmed-medical-dataset-2014-to-2023-title-abstract containing 142,432 papers from 2014 to 2023 in the biomedical field. These diverse datasets ensure robust coverage of both general and domain-specific scholarly topics for analysis.
@@ IV-B Experiment Setup
Baseline Methods. We evaluate our model against ten benchmark models encompassing a spectrum of traditional and advanced neural topic modeling techniques: (i) LDA [1], a classical probabilistic generative model for topic discovery. (ii) NMF [2], a matrix decomposition technique used for topic discovery. (iii) ProdLDA [38], a neural topic model based on variational autoencoders, incorporating product-of-experts priors for topic generation. (iv) DecTM [39] decouples topic modeling into separate modules for word and document distributions. (v) ETM [7] combines word embeddings with generative topic modeling. (vi) NSTM [11] utilizes optimal transport theory to mitigate semantic bias in neural topic models. (vii) TSCTM [13] employs contrastive learning for short text topic modeling. (viii) ECRTM [15] prevents topic collapse through embedding clustering regularization. (ix) BERTopic [18] leverages pre-trained transformer-based embeddings for topic generation. (x) FASTopic [16] introduces dual semantic-relation reconstruction for adaptive, stable, and transferable topic discovery. We fine-tune the hyperparameters of these baselines under different datasets and topic numbers.
Implementation Details. We use the BGE-M3 model for embedding generation, fine-tuned on domain-specific data to optimize semantic representation [40]. Clustering uses the K-Means algorithm [41] and the Llama-3.1-70B model [42]. Parameters are set as $\alpha=1$ and $\lambda=0.5$, and experiments are conducted on two NVIDIA A100-80GB GPUs.
@@ IV-D Case Study
We present three illustrative case studies to thoroughly assess the effectiveness and interpretability of SciTopic: the first focuses on clustering accuracy using ground-truth labels, the second examines source venues of papers associated with each topic, and the third investigates topics’ temporal evolution over time.
@@
To assess topic interpretability over time, we analyze the dynamics of Topic 2 (computer vision) and Topic 5 (information retrieval), as shown in Figure 4. Topic 2 shows marked growth after 2013, which reflects the rapid advancement of CNNs in the computer vision field after the breakthrough of AlexNet in 2012. And Topic 5 shows a clear peak around 2008-2012, followed by a gradual decline. This trend was likely associated with peak research interest in information retrieval during the rapid expansion of search engines such as Bing and Baidu, as well as ongoing innovation at Google.

### v2

@@ Abstract
Topic discovery in scientific literature provides valuable insights for researchers to identify emerging trends and explore new avenues for investigation, facilitating easier scientific information retrieval. Many machine learning methods, particularly deep embedding techniques, have been applied to discover research topics. However, most existing topic discovery methods rely on word embedding to capture the semantics and lack a comprehensive understanding of scientific publications, struggling with complex, high-dimensional text relationships. Inspired by the exceptional comprehension of textual information by large language models (LLMs), we propose an advanced topic discovery method enhanced by LLMs to improve scientific topic identification, namely SciTopic. Specifically, we first build a textual encoder to capture the content from scientific publications, including metadata, title, and abstract. Next, we construct a space optimization module that integrates entropy-based sampling and triplet tasks guided by LLMs, enhancing the focus on thematic relevance and contextual intricacies between ambiguous instances. Then, we propose to fine-tune the textual encoder based on the guidance from the LLMs by optimizing the contrastive loss of the triplets, forcing the text encoder to better discriminate instances of different topics. Finally, extensive experiments conducted on three real-world datasets of scientific publications demonstrate that SciTopic outperforms the state-of-the-art (SOTA) scientific topic discovery methods, enabling researchers to gain deeper and faster insights [^1] Access the source code link: https://github.com/CNICDS/SciTopic.
@@ I Introduction
Recent advancements in machine learning, particularly in the realm of deep learning [xu2025scsiameseclu, wang2024deep, zhang2025motif], have led to the emergence of various techniques aimed at automating the topic discovery process. The classical topic modeling techniques, including Latent Dirichlet Allocation (LDA) [blei2003latent], Non-negative Matrix Factorization (NMF) [lee2000algorithms], and Probabilistic Latent Semantic Analysis (PLSA) [hofmann1999probabilistic], could be applied to discover the scientific topics directly. However, these bag-of-words methods ignore contextual word interrelations, failing to capture the intricate semantics of modern scientific texts. In addition, these techniques often necessitate dimensionality reduction processes such as Principal Component Analysis (PCA) [abdi2010principal] or Uniform Manifold Approximation and Projection (UMAP) [mcinnes2018umap], potentially leading to significant information loss that is vital for maintaining the thematic depth of the documents. Deep embedding methods have gained prominence for their ability to represent textual data in a high-dimensional space, capturing semantic relationships between words and phrases. Unlike traditional bag-of-words approaches that treat words in isolation, document embeddings encapsulate the overall significance of a document in a continuous vector space, aligning semantically akin words more closely [mikolov2013efficient]. Even advanced deep topic modeling like the Embedded Topic Model (ETM) [dieng2020topic] and Neural Variational Document Model (NVDM) [miao2016neural], though incorporating word embeddings to capture semantic nuances, still results in a limited understanding of the intricate relationships. These limitations ultimately impact the quality of insights derived from topic discovery, potentially leading to incomplete or inaccurate representations of the underlying thematic structures within the literature.
@@
To overcome these limitations, we draw inspiration from the remarkable capabilities of large language models (LLMs) in comprehending textual information. LLMs, such as GPT-4 and BERT, have exhibited unparalleled proficiency in natural language comprehension by recognizing deep contextual connections between words, phrases, and entire texts. Transformer-based architectures can capture long-range dependencies and process substantial volumes of text in a context-sensitive manner [vaswani2017attention], effectively discerning nuanced thematic structures. Trained on vast data, transformer-based models can capture complex patterns, contextual cues, and semantic relationships beyond mere word co-occurrence. These embeddings retain syntactic and semantic linkages, offering comprehensive depictions of documents’ thematic content compared to earlier sparse models. This enhancement allows for refined topic modeling and improved semantic clustering, naturally grouping documents based on their intrinsic meanings rather than mere word frequency. For instance, Sentence-BERT (SBERT) optimizes BERT embeddings for semantic similarity tasks using a siamese network structure, preserving intricate semantic details and capturing subtle thematic distinctions [reimers2019sentence].
@@
Along this line, by leveraging the strengths of LLMs, we propose SciTopic, an effective method to enhance scientific topic identification and provide deeper insights for researchers. Firstly, we construct a text encoder that captures essential content from scientific publications, including metadata, titles, and abstracts. Through this module, we can extract meaningful textual features crucial for accurate topic identification. Then, we introduce an LLM-guided clustering technique that leverages entropy-based sampling and triplet tasks. This approach diverges from traditional unsupervised clustering methods by actively involving the LLM in the clustering process. We utilize an entropy-based sampling strategy to identify the most ambiguous or uncertain documents, where cluster membership is less defined. These high-entropy instances are used as anchors for the triplet tasks, where two candidate titles or abstracts from nearby clusters are selected. By analyzing these triplets, the LLM refines document embeddings, sharpening the distinctions between closely related clusters. This method not only improves clustering precision but also minimizes computational overhead by focusing the LLM’s attention toward the most informative cases. Ultimately, this approach leads to a more contextually accurate and thematically coherent clustering result, ensuring that even subtle topic differences are effectively captured. Our key contributions are as follows:
@@ II-A Neural Topic Models
Neural Topic Models (NTMs) integrate deep learning with probabilistic modeling to improve topic coherence and document representation [lee2000algorithms, zhao2020neural, dieng2020topic, Zhao2021TopicMM, wu2022mitigating, yang-etal-2025-neural]. LDA inspires many neural adaptations addressing scalability and coherence for large vocabularies [blei2003latent]. Recent advancements include ECRTM, which prevents topic collapse through embedding clustering regularization [wu2023effective], and FASTopic, utilizing dual semantic-relation reconstruction to regulate embeddings [wu2024fastopic]. InfoCTM aligns topics cross-lingually using mutual information [wu2023infoctm], while BERTopic leverages pre-trained transformers and clustering for coherent topics [grootendorst2022bertopic]. Despite these advancements, most NTMs are not tailored for scientific topic discovery tasks.
@@ II-B Scientific Topic Discovery
Scientific topic discovery automates the identification of research trends and emerging areas, essential given the rapid growth in publications [langley1987scientific, Huang2022TopicDI, ninkov2022bibliometrics]. Domain-specific models assess research impact [Guo2014ATT, Zhang2018DomainspecificTM]. For instance,  [PonsPorrata2007TopicDB] adapts hierarchical topic discovery for scientific literature, and  [Jeong2016DiscoveryOR] tracks authors’ research evolution.  [Altaf2021ScientificDD] proposes dataset recommendation at the topic level, while  [Wu2024EmergingST] identifies emerging topics by analyzing rare synonymous biterms. However, most methods fail to adequately capture complex relationships.
@@ II-C Prompt Learning
Prompt learning advances LLM applications by designing task-specific cues, enabling few-shot and zero-shot scenarios [Liu2021PretrainPA, He2023YouOP]. Techniques like domain-controlled prompts for remote sensing [Cao2023DomainControlledPL], graph prompt learning [Sun2023GraphPL], and Match-Prompt frameworks for diverse tasks [Xu2022MatchPromptIM] have demonstrated effectiveness. Methods such as region-based image recognition [Zhou2023HeterogeneousRE], unsupervised image enhancement [Liang2023IterativePL], and continual learning [Tang2023WhenPI] further showcase prompt learning’s versatility. Inspired by prompt-based learning, we use prompts to improve query triplet generation for scientific topic discovery.
@@ III-A Problem Definition
In our task, we aim to extract meaningful research topics from large collections of scientific papers. Each paper is in the form of a document containing title ($t$), abstract ($a$), and metadata ($m$) and can be represented by a set of textual features $\{x_{1},x_{2},...,x_{n}\}$}. Then, we cluster documents into distinct groups $\{C_{1},C_{2},...,C_{k}\}$}. From each cluster, we further extract a list of key terms to verbalize the represented topic, enabling efficient topic identification and trend analysis.
@@ III-B Framework Overview
As depicted in Figure 1, our method starts with encoding the title, abstract, and metadata of each document and concatenating their embeddings to a composite feature matrix. Then we group these documents into clusters and sample ambiguous instances to construct the triplet task, which prompts the LLM to evaluate and reassign documents to more coherent clusters. These refined triplets are finally used with contrastive learning [ning2025rethinkinggraphcontrastivelearning] to fine-tune the embedding model for an enhanced clustering performance.
@@ III-C Textual Encoder
Each document is divided into three key components: the title, abstract, and metadata. The title and abstract are encoded separately using a fine-tuned model capable of handling varied text granularities. Metadata is consolidated into a single string and embedded using the same model to maintain consistency. The encoding process is expressed as follows:
@@
III-D LLM-Guided Clustering
Here, we refine text clustering using a triplet task guided by large language models (LLMs), designed to reflect user-specific perspectives. Each triplet comprises an anchor and two candidate elements $(a,c_{1},c_{2})$, aiming to identify which candidate is more thematically similar to the anchor.
@@ III-E Fine-tuning
The fine-tuning process employs a cross-entropy loss with in-batch negative sampling [wang2021cross]:
@@
where $f(x)$ is embedding function, and $\gamma$ is a margin parameter that helps avoid trivial solutions by enforcing a minimal semantic distinction between positive and negative pairs.
@@ III-F Topic Verbalization
As depicted in Figure 2, we use class-based term frequency-inverse document frequency (c-TF-IDF) [grootendorst2022bertopic], which calculates scores for each cluster by treating it as a single document, emphasizing key terms that define each topic. The weight of a term $t$ in cluster $c$ is calculated as:
@@ IV-A Datasets
This study identifies prevalent topics in scholarly articles using the DBLP database and two additional Kaggle datasets, as shown in Table I. The AI-DM Research Literature Dataset, based on the DBLP database, includes metadata for 57,320 AI papers and 20,700 DM papers from major conferences such as AAAI, ACL, CVPR, IJCAI, NeurIPS, and SIGKDD, covering titles, authors, publication years, and venues. Additionally, the DBLP V10222https://www.kaggle.com/datasets/nechbamohammed/research-papers-dataset Dataset, derived from DBLP (Version 10), spans fields like computer science, mathematics, and physics, offering metadata for 999,064 papers, including 827,533 with both titles and abstracts, from which 100,000 papers were randomly sampled. The NeurIPS [^3] https://www.kaggle.com/datasets/benhamner/nips-papers Dataset provides detailed information on 7,241 NeurIPS conference papers from 1987 to 2016, including titles, authors, abstracts, and full texts, offering comprehensive insights into machine learning advancements. This combination of datasets ensures robust coverage of both general and domain-specific scholarly topics for analysis. Furthermore, this study incorporates additional datasets for broader coverage, including 20,000 randomly sampled papers from the arXiv [^4] https://www.kaggle.com/datasets/Cornell-University/arxiv, and the PubMed Dataset [^5] https://www.kaggle.com/datasets/nabarupghosh/pubmed-medical-dataset-2014-to-2023-title-abstract containing 142,432 papers from 2014 to 2023 in the biomedical field. These diverse datasets ensure robust coverage of both general and domain-specific scholarly topics for analysis.
@@ IV-B Experiment Setup
Baseline Methods. We evaluate our model against ten benchmark models encompassing a spectrum of traditional and advanced neural topic modeling techniques: (i) LDA [blei2003latent], a classical probabilistic generative model for topic discovery. (ii) NMF [lee2000algorithms], a matrix decomposition technique used for topic discovery. (iii) ProdLDA [srivastava2017autoencoding], a neural topic model based on variational autoencoders, incorporating product-of-experts priors for topic generation. (iv) DecTM [wu2021discovering] decouples topic modeling into separate modules for word and document distributions. (v) ETM [dieng2020topic] combines word embeddings with generative topic modeling. (vi) NSTM [zhao2020neural] utilizes optimal transport theory to mitigate semantic bias in neural topic models. (vii) TSCTM [wu2022mitigating] employs contrastive learning for short text topic modeling. (viii) ECRTM [wu2023effective] prevents topic collapse through embedding clustering regularization. (ix) BERTopic [grootendorst2022bertopic] leverages pre-trained transformer-based embeddings for topic generation. (x) FASTopic [wu2024fastopic] introduces dual semantic-relation reconstruction for adaptive, stable, and transferable topic discovery. We fine-tune the hyperparameters of these baselines under different datasets and topic numbers.
Implementation Details. We use the BGE-M3 model for embedding generation, fine-tuned on domain-specific data to optimize semantic representation [chen2024bge]. Clustering uses the K-Means algorithm [macqueen1967some] and the Llama-3.1-70B model [dubey2024llama]. Parameters are set as $\alpha=1$ and $\lambda=0.5$, and experiments are conducted on two NVIDIA A100-80GB GPUs.
@@ IV-D Case Study
We present three case studies to thoroughly assess the effectiveness and interpretability of SciTopic: the first focuses on clustering accuracy using ground-truth labels, the second examines source venues of papers associated with each topic, and the third investigates topics’ temporal evolution over time.
@@
To assess topic interpretability over time, we analyze the dynamics of Topic 2 (computer vision) and Topic 5 (information retrieval), as shown in Figure 4. Topic 2 shows marked growth after 2013, which reflects the rapid advancement of CNNs in the computer vision field after the breakthrough of AlexNet in 2012. And Topic 5 shows a clear peak around 2008-2012, followed by a gradual decline. This trend was likely associated with peak research interest in information retrieval during the rapid expansion of search engines such as Google, Bing and Baidu.